# On the Effectiveness of Methods and Metrics for Explainable AI in Remote Sensing Image Scene Classification

### v1

@@ II B 1 Background
II B 2 Limitations
Gradient-propagation-based methods exhibit limitations when applied to RS imagery, particularly when multiple LULC classes occur in a scene. For instance, in a RS image containing five airplanes, GradCAM localizes only a single instance [42], underscoring its inability to resolve multiple objects of interest at varying scales [43, 44]. This failure originates from the reliance of GradCAM on a global gradient averaging mechanism, which collapses spatial gradient information into a single channel-wise importance weight. For two distinct objects $O_{1}$ and $O_{2}$, activating disjoint regions $R_{1}$ and $R_{2}$ in $A^{p}$, the gradients $\nabla_{1}$ and $\nabla_{2}$ are:
@@ II B 2 Limitations
When $\nabla_{1}\neq\nabla_{2}$, the larger gradient contributes more to the sum, so the computed weight is more reflective of the larger gradient. However, because $\omega_{p}^{c}$ is a scalar applied uniformly across the feature map, the distinct spatial contributions from $R_{1}$ and $R_{2}$ are merged. Consequently, the resulting attribution map $\Phi_{\mathrm{GradCAM}}^{c}$ loses the spatial granularity necessary to differentiate between the contributions of multiple objects.
II C Relevancy-propagation-based methods
II C 1 Background
Relevancy-propagation-based methods, such as LRP or DeepLIFT, explain the predictions of classifiers by back-propagating a relevancy score throughout the network, to attribute importance to each input feature. For relevancy-propagation-based explanation methods, we consider DeepLIFT [12] and LRP [11].
DeepLIFT [12] attributes the difference in the outputs $f_{c}(\bm{x})-f_{c}\left(\tilde{\bm{x}}\right)$ on the differences between the input image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ and a perturbation baseline (reference) $\tilde{\bm{x}}\;=\;\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$. Specifically, DeepLIFT runs forward propagation for both $\bm{x}$ and $\tilde{\bm{x}}$ inputs. Then, it stores the difference of the weighted activation $\Delta z_{s,t}=w_{s,t}^{(l,l+1)}\left(\bm{x}_{s}^{l}-\tilde{\bm{x}}^{l}_{s}\right)$, where $\bm{x}_{s}^{l}$ and $\tilde{\bm{x}}^{l}_{s}$ are the activations of the neuron $s$ in the $l$ layer and $w_{s,t}^{(l,l+1)}$ is the weight between neuron $s$ and neuron $t$ in the next layer $l+1$. Finally, the backpropagation of the difference in the outputs to the final input contributions $\Phi_{\text{DeepLift}}$ is achieved with the following rules:
@@ II C 1 Background
and finally
@@
Layer Relevance Propagation (LRP)[11] assigns a relevance score $r$ to each input feature of a network by propagating the output of the model back through the network to the input layer. LRP interprets the neural network as a flow graph, redistributing prediction relevance backwards through the layers using specific propagation rules.
Relevance scores $r$ for a given layer $l$ are propagated to the neurons in the preceding layer as follows:
@@
where $z_{s,t}$ represents the contribution of neuron $s$ in layer $l$ to the activation of neuron $t$ in layer $l+1$, ensuring that total relevance is conserved across layers. LRP employs specific propagation rules to distribute relevance back to input features depending on the network architecture and task; see Montavon et al. [45] for details. For the final layer, the relevance $r_{t}^{(L)}$ is initialized as the output score for the target class $c$. The LRP method $\Phi_{\text{LRP}}$ for attributing the input space is defined as:
@@
Figure 3: LRP visualizations for an image from the DG-ML dataset. a) Original image showing two classes: Urban Land and agricultural Land; b) Pixel-wise reference map (blue: Urban Land; orange: agricultural Land); c) LRP explanation for agricultural Land (normalized range: 0–1, unnormalized range: 0–0.02); d) LRP explanation for Urban Land (normalized range: 0–1, unnormalized range: 0–0.18).
II C 2 Limitations
As DeepLIFT calculates relevance by comparing the output of a model for the input $\bm{x}$ and a reference input $\tilde{\bm{x}}$, expressed as $f_{c}(\bm{x})-f_{c}\left(\tilde{\bm{x}}\right)$, it encounters similar issues to perturbation-based methods (see Section  II A 2). Sturmfels et al. [46] investigate the impact of various reference baselines for the Integrated Gradients method, which also employs a difference-from-reference approach analogous to DeepLIFT. Their study visually examines reference baselines such as Gaussian blur, uniform values, blurring, maximum distance, and black baselines. They find that the selection of an appropriate baseline requires domain-specific knowledge to ensure it accurately represents missingness within the data domain. This challenge is amplified in RS image scene classification due to the availability of various spectral bands in the data.
A central property of LRP and related relevancy-propagation-based methods such as Integrated Gradients [47] is relevance conservation: the total relevance assigned to the input for class $c$ equals the network output, $\sum_{j}r_{j}^{c}=f_{c}(\bm{x})$ [45]. In natural images, discriminative information is typically concentrated within the object of interest. In contrast, RS scenes usually contain several LULC classes, so that a single class may occur in multiple, spatially disconnected regions; the discriminative information is therefore distributed rather than localized. Because the overall relevance budget is fixed, two classes that yield comparable output scores but occupy differently sized areas receive different relevance densities: the larger region is attributed a lower per-pixel relevance than the smaller region. Fig. 3 illustrates this issue using an image from the DG-ML dataset containing differently sized two classes: agricultural land ($c_{\text{major}}$) and urban land ($c_{\text{minor}}$). As the outputs of the model for both classes are similar $(f_{c_{\text{major}}}(\bm{x})\approxeq f_{c_{\text{minor}}}(\bm{x}))$ the maximum relevance to be propagated by LRP is similar. The higher number of relevant pixels for $c_{\text{major}}$ results in lower attributed relevance per pixel compared to $c_{\text{minor}}$. The $r^{c}_{j}$ for $c_{\text{major}}$ span values in the interval $[0,0.02]$, while the $r^{c}_{j}$ for $c_{\text{minor}}$ range within $[0,0.18]$. As attributions are typically normalized between 00 and $1$, this can produce attribution maps that are highly sensitive to outliers.
@@ III Explanation metrics for RS image scene classification
In this section, we first provide a brief overview of the categorization of explanation metrics. We then introduce selected explanation metrics from different categories and discuss their suitability to estimate the quality of explanations for RS image scene classification. Explanation metrics offer a quantitative way to assess the quality of explanation methods. An explanation can be defined as a “presentation […] in human-understandable terms” [22]. Because this definition centers on subjective human interpretation, it inherently lacks a single, absolute ground truth. Therefore, it is not possible to estimate an explanation error, as one would estimate the performance error of a DL model [24]. Furthermore, explanation quality depends on more than the explanation method itself. When the decision-making process of a model is erroneous, even correctly derived explanations can appear inconsistent or counterintuitive, since explanations inherently reflect the internal logic of the model [48]. To address these challenges, researchers have proposed metrics across several categories that capture distinct aspects of explanation quality. In our study, we consider five different categories of explanation metrics: Faithfulness, Robustness, Localization, Complexity and Randomization. Formally, the quality of an explanation is assessed using a metric $\Psi$, which evaluates the effectiveness of an explanation method:
@@
where $\hat{q}$ is a scalar quality score. For simplicity, we omit explicit parameterization unless necessary.
III A Faithfulness
III A 1 Background
The first category that we consider is faithfulness, which measures how well explanations reflect the actual decision-making process of the model. A faithful explanation accurately reflects why the model made a specific prediction. Metrics in this category assess this alignment by evaluating whether features identified as important, significantly influence the prediction of the model. Let us consider $\bm{x}$ as the input image and $V\subseteq[H]\times[W]$ a set of pixels to remove. We define the perturbation function $\mathcal{P}_{\mathbb{X}}(\bm{x},V)$ which replaces each feature in $V$ with a predefined baseline value. The general faithfulness category $\Psi_{\mathrm{F}}(\bm{x},V)$ is defined as:
$\Psi_{\mathrm{F}}(\bm{x},V)=\bigl{\lvert}\,f(\bm{x})-f\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},V)\bigr{)}\bigr{\rvert}.$ (22)
Perturbations are commonly applied to the most relevant pixels first (MoRF) or the least relevant ones first (LeRF).
To evaluate explanation methods for RS image scene classification, we consider two faithfulness metrics: i) Faithfulness Estimate (FE)[28], which measures the correlation between feature importance and prediction change, and ii) Iterative Removal Of Features (IROF)[29], which assesses prediction certainty decay when removing the most relevant segments.
Faithfulness Estimate (FE)[28] quantifies the alignment between attribution magnitude and model sensitivity by measuring the correlation between the total attribution on the top-$K$ pixels and the resulting drop in prediction when those pixels are removed. Formally, we define the FE metric $\Psi_{\mathrm{FE}}$ as:
@@ III A 1 Background
Here, $\Phi(\bm{x})_{j,k}$ denotes the attributed relevancy at pixel $(j,k)$, and $V\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ is the set of the $K$ pixel indices with the largest attribution values. The function $\mathrm{corr}$ denotes the Pearson correlation coefficient. To compute $\Psi_{\mathrm{FE}}$, one first ranks all pixels by $\Phi(\bm{x})_{j,k}$, selects the top-$K$ indices $V$, and then correlates the summed attributions on $V$ with the change $f(\bm{x})-f\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},V)\bigr{)}$. A higher value of $\Psi_{\mathrm{FE}}$ indicates that the explanation method attributes the behavior of the model, i.e., i.e., removing features deemed important by the method results in a significant drop in prediction certainty.
Iterative Removal Of Features (IROF)[29] measures how the prediction certainty changes when the most relevant features are removed but it considers the $K$ most relevant segments rather than pixels. Let us define $\{S_{m}\}_{m=1}^{M}$ as a partition of the image $\bm{x}$ into $M$ disjoint segments and sort them by their relevance, that:
$\frac{1}{|S_{1}|}\sum_{(j,k)\in S_{1}}\Phi(\bm{x})_{j,k}\;\geq\;\cdots\;\geq\;\frac{1}{|S_{M}|}\sum_{(j,k)\in S_{M}}\Phi(\bm{x})_{j,k}.$ (24)
For each $\kappa=0,1,\dots,K$, define the cumulative removal set $V_{\kappa}\;=\;\bigcup_{m=1}^{\kappa}S_{m},$ and the corresponding perturbed input $\mathcal{P}_{\mathbb{X}}(\bm{x},V_{\kappa}).$ Finally, the IROF score $\Psi_{\mathrm{IROF}}$ for $\bm{x}$ is obtained by normalizing each perturbed model output by the original confidence:
@@
III A 2 Limitations
Faithfulness metrics inherit several limitations from perturbation-based explanation methods, as both rely on input perturbations to evaluate changes in the prediction certainty of a model. Consequently, they exhibit similar issues to those discussed in Section II A 2. For example, like perturbation-based methods, most faithfulness metrics require a carefully chosen baseline to avoid distribution shifts that can affect the predictions of the model.
@@ III A 2 Limitations
In addition, the FE metric relies on a strong assumption: that removing a small number of high-attributed pixels will lead to a measurable change in model output. This assumption implicitly presumes that class-relevant information is concentrated in a small subset of pixels. However, RS scenes often contain multiple spatially disjoint regions associated with the same LULC class. These regions exhibit redundancy, as similar class-relevant features are repeated across the scene. In such cases, removing a small number of pixels does not produce a meaningful change in prediction, and the attribution sum remains nearly constant across different pixel subsets. As a result, the Pearson correlation used by FE may become numerically unstable, since both numerator and denominator can approach zero, leading to undefined or misleading scores even for faithful explanations.
In contrast to FE, some faithfulness metrics, such as IROF, evaluate the behavior of the model over a sequence of increasingly perturbed inputs. For these trajectory-based metrics, the order in which features are removed plays a critical role in shaping the evaluation outcome. To illustrate this, let us consider a scene in which a class $c_{\text{major}}$ occupies most of the image area, whereas a different class $c_{\text{minor}}$ covers only a small region. In such images, perturbing pixels belonging to $c_{\text{major}}$ do not significantly impact prediction certainty, as multiple pixels or regions contain similar predictive information. Consequently, metrics employing the MoRF strategy, which sequentially removes the most relevant pixels first, fail to capture meaningful changes in prediction certainty, as demonstrated by the slow decline of $f(\bm{x})$ in Fig. 4. In contrast, perturbations of the fewer, highly informative pixels associated with $c_{\text{minor}}$ lead to a rapid decrease in prediction certainty. This discrepancy results in biased metric outcomes, incorrectly suggesting that explanations for $c_{\text{minor}}$ are faithful, while explanations for $c_{\text{major}}$ appear less faithful. The LeRF strategy, which removes the least relevant pixels first, provides a better alternative to address this issue. By preserving the crucial predictive features of $c_{\text{minor}}$ until later perturbation steps, the LeRF strategy ensures a stable prediction certainty and yields a more balanced evaluation of explanation methods.
@@
Figure 4: Comparison of MoRF and LeRF removal strategies for the target class $c$ (agricultural land, orange). Here, $K$ is the percentage of removed pixels and $f_{c}(\bm{x})$ represents the prediction certainty. Top row: a) Original image from the DG-ML dataset; b) Pixel-wise reference map. Agriculture land in orange; MoRF removal: c) $K=10\%$, $f_{c}(\bm{x})=0.98$; d) $K=50\%$, $f_{c}(\bm{x})=0.64$; e) $K=90\%$, $f_{c}(\bm{x})=0.31$. Bottom row: f) Grad-CAM attribution map for $c$; LeRF removal: g) $K=10\%$, $f_{c}(\bm{x})=0.99$; h) $K=50\%$, $f_{c}(\bm{x})=0.99$; i) $K=90\%$, $f_{c}(\bm{x})=0.72$.
III B Robustness
III B 1 Background
The second category we consider is robustness, which assesses the stability of explanations under small input perturbations. Similar, to adversarial attacks on DL models [49], explanations can also be fooled [50, 51]. Thus, ensuring that the explanation is invariant to minimal perturbations in the input space is crucial for critical applications. We require small perturbations $\delta$ such that $\left\|\delta\right\|_{2}<\varepsilon$ and expect $\left\|\mathcal{P}_{\mathbb{X}}(\bm{x})-\bm{x}\right\|_{2}<\varepsilon$ for a small positive constant $\varepsilon$, assuming that the model output approximately stayed the same $f(\bm{x})\approx f\left(\mathcal{P}_{\mathbb{X}}(\bm{x})\right)$. The category $\Psi_{\mathrm{ROB}}$ can be defined as:
$\Psi_{\mathrm{ROB}}(\Phi,\bm{x},\varepsilon)=\max_{\|\delta\|_{2}\leq\varepsilon}\bigl{\|}\,\Phi(\bm{x})\;-\;\Phi(\bm{x}+\delta)\bigr{\|}_{2}.$ (26)
@@ III B 1 Background
To evaluate the robustness of explanation methods for RS image scene classification, we use i) Average Sensitivity (AS)[30], which measures sensitivity to small perturbations using a Monte Carlo-based approximation and ii) Local Lipschitz Estimate (LLE)[28], which estimates Lipschitz continuity to assess explanation stability under minor input changes.
@@
Average Sensitivity (AS)[30] evaluates the sensitivity of the explanation method by perturbing samples using a Monte-Carlo-based approximation. The AS metric $\Psi_{\mathrm{AS}}$ is defined as:
@@
where $\varepsilon$ defines the radius of a discrete, finite-sample neighborhood around each input sample $\bm{x}$. This neighborhood denoted as $\mathcal{N}_{\epsilon}(\bm{x})$, includes all samples that are within a distance of $\varepsilon$ from $\bm{x}$. A lower score is indicative of more robustness.
The Local Lipschitz Estimate (LLE)[28] estimates the Lipschitz continuity of the explanation, which measures how much the explanation changes with respect to the input under a minor perturbation, meaning the prediction did not change. The LLE metric $\Psi_{\mathrm{LLE}}$ is defined as:
@@
where lower values indicate less change concerning the change in input, which is desirable.
III B 2 Limitations
A key hyper-parameter in this metric category is the noise type. If the introduced noise does not naturally occur in the acquired images, it may generate OoD samples, reducing the reliability of the model and thereby affecting the validity of the explanations. When evaluating the robustness of explanation methods, it is important to introduce noise that does not result in OoD samples. This consideration also depends on the model architecture, as many models exhibit inherent resilience to certain types of noise. For instance, in SAR images, multiplicative noise such as speckle is pervasive [52]. In contrast, for multispectral images, noise is typically independent of the data and generally additive in nature, such as additive white Gaussian noise [53, 54].
III C Localization
III C 1 Background
The third category we consider is localization, which measures the spatial accuracy of explanations by comparing them to reference maps or bounding boxes. This comparison can be quantified using a similarity measure SIM between the generated explanation map $\Phi(\bm{x})$ and a pixel-level reference map $\bm{s}_{gt}$. The category $\Psi_{\mathrm{LOC}}$ is defined as:
$\Psi_{\mathrm{LOC}}(\Phi,\bm{x},\bm{s}_{gt},\text{SIM})=\text{SIM}\left(\Phi(\bm{x}),\bm{s}_{gt}\right),$ (29)
where SIM can be a measure such as Intersection over Union, Pearson correlation, or SSIM depending on the nature of the task and reference map.
To evaluate the localization capabilities of explanation methods for RS image scene classification, we use i) Top-K Intersection (TKI)[32], which measures the fraction of top-ranked attributions within the pixel-level reference map, and ii) Relevance Rank Accuracy (RRA)[31], which assesses the concentration of high-relevance values within the reference map.
Top-K Intersection (TKI), as proposed by Theiner et al. [32], measures the fraction of the $K$ highest ranked attributions that lie on the binary reference mask $\bm{s}_{gt}$. Let us define
$I_{K}=\{\,i_{1},i_{2},\ldots,i_{K}\}\quad\text{and}\quad S_{gt}=\{\,p\mid\bm{s}_{gt}(p)=1\},$ (30)
where $I_{K}$ is the set of the top-$K$ ranked pixel indices and $S_{gt}$ the set of true-positive pixels in the mask. The intersection of these sets is:
$\bm{s}_{1-K}=I_{K}\;\cap\;S_{gt}.$ (31)
The TKI metric $\Psi_{\mathrm{TKI}}$ is then defined as:

### v3

@@ II.B.1 Background
II.B.2 Limitations
Gradient-propagation-based methods exhibit limitations when applied to RS scenes with multiple LULC classes. For instance, in a scene containing five airplanes, GradCAM localizes only a single instance [47], underscoring its inability to resolve multiple objects of interest at varying scales [48, 49]. This failure originates from the reliance of GradCAM on a global gradient averaging mechanism, which collapses spatial gradient information into a single channel-wise importance weight. For two distinct objects $O_{1}$ and $O_{2}$, activating disjoint regions $R_{1}$ and $R_{2}$ in $A^{p}$, the gradients $\nabla_{1}$ and $\nabla_{2}$ are:
@@ II.B.2 Limitations
When $\nabla_{1}\neq\nabla_{2}$, the larger gradient contributes more to the sum, so the computed weight is more reflective of the larger gradient. Since $\omega_{p}^{c}$ is a scalar applied uniformly to $A^{p}$, spatially distinct contributions from $R_{1}$ and $R_{2}$ merge, and the attribution map $\Phi_{\mathrm{GradCAM}}^{c}$ loses the granularity needed to separate multiple objects.
II.C Relevancy-propagation-based methods
II.C.1 Background
Relevancy-propagation-based methods explain predictions by back-propagating relevance scores to attribute importance to input features. We consider DeepLIFT [12] and LRP [11].
[…]
@@ II.C.1 Background
and the final attribution $\Phi_{\text{DeepLift}}$ is defined as:
@@
LRP [11] propagates output relevance back to inputs by interpreting the network as a flow graph. For a layer $l$, relevance scores $r$ are redistributed as follows:
@@
where $z_{s,t}$ is the contribution of neuron $s$ in layer $l$ to neuron $t$ in $l+1$, ensuring conservation of total relevance. Specific propagation rules depend on the model architecture and task [50]. For the final layer, $r_{t}^{(L)}$ is initialized as the target class score $f_{c}(\bm{x})$. The input attribution LRP method $\Phi_{\text{LRP}}$ is then:
@@
Figure 3: LRP visualizations for an image from the DG-ML [46] dataset. a) Original image with two classes: Urban Land and Agricultural Land; b) Pixel-wise reference map (blue: Urban Land, orange: Agricultural Land); c) LRP explanation for Agricultural Land (normalized range: [0,1], unnormalized range: [0,0.02]); d) LRP explanation for Urban Land (normalized range: [0,1], unnormalized range: [0,0.18]).
II.C.2 Limitations
[…]
A central property of LRP and related methods such as Integrated Gradients [52] is relevance conservation: the total input relevance equals the network output, $\sum_{j}r_{j}^{c}=f_{c}(\bm{x})$ [50]. RS scenes usually contain several LULC classes, so that a single class may occur in multiple, spatially disconnected regions. The discriminative information is therefore distributed rather than localized. Because the overall relevance budget is fixed, two classes that yield comparable output scores but occupy differently sized areas receive different relevance densities. The larger region is attributed a lower per-pixel relevance than the smaller region. Fig. 3 shows this discrepancy for a DG-ML image containing agricultural land ($c_{\text{major}}$) and urban land ($c_{\text{minor}}$). Since $f_{c_{\text{major}}}(\bm{x})\approxeq f_{c_{\text{minor}}}(\bm{x})$, both propagate similar total relevance. However, the larger $c_{\text{major}}$ region yields lower per-pixel relevance ($[0,0.02]$) than $c_{\text{minor}}$ ($[0,0.18]$). As attributions are typically normalized to $[0,1]$, this discrepancy increases sensitivity to outliers.
@@ III Explanation metrics for RS image scene classification
In this section, we first provide a brief overview of the categorization of explanation metrics. We then introduce selected explanation metrics from different categories and discuss their suitability to estimate the quality of explanations for RS image scene classification. Explanation metrics quantify the quality of explanation methods. An explanation is commonly defined as a “presentation […] in human-understandable terms” [22]. Since this definition depends on subjective interpretation, there is no absolute ground truth and thus no direct notion of explanation error [24]. Furthermore, explanation quality also depends on the model itself: if the model’s reasoning is flawed, even faithful explanations may appear inconsistent or counterintuitive [53].
To address these challenges, different metrics have been proposed across different categories, each capturing distinct aspects of explanation quality. We consider five categories of explanation metrics: Faithfulness, Robustness, Localization, Complexity, and Randomization. Formally, explanation quality is measured by a metric $\Psi$ evaluating the effectiveness of a method $\Phi$:
@@
where $\hat{q}$ is a scalar quality score. Explicit parameterization is omitted unless required.
III.A Faithfulness
III.A.1 Background
Faithfulness measures how well explanations reflect the decision process of the model. A faithful explanation identifies features that truly influence predictions. Given input $\bm{x}$ and a set of pixels $V\subseteq[H]\times[W]$ to remove, the perturbation function $\mathcal{P}_{\mathbb{X}}(\bm{x},V)$ replaces the pixel in $V$ with a baseline value. Perturbations are applied either to the most relevant features first (MoRF), the least relevant first (LeRF), or randomly. We consider two faithfulness metrics: i) Faithfulness Estimate (FE) [28], which measures the correlation between feature importance and prediction change, and ii) Iterative Removal Of Features (IROF) [29], which evaluates prediction certainty decay when removing relevant segments.
FE [28] measures the correlation between summed attributions on top-$K$ pixels and the prediction drop when those pixels are removed. Formally, we define the FE metric $\Psi_{\mathrm{FE}}$ as:
@@ III.A.1 Background
Here, $\Phi(\bm{x})_{j,k}$ denotes the attributed relevancy at pixel $(j,k)$, and $V\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ is the set of the $K$ pixel indices with the largest attribution values and $\mathrm{corr}$ denotes the Pearson correlation coefficient. A high $\Psi_{\mathrm{FE}}$ value indicates that removing highly attributed pixels substantially lowers the prediction certainty.
IROF [29] evaluates prediction certainty decay when removing the $K$ most relevant segments rather than pixels. Let $\{S_{m}\}_{m=1}^{M}$ be a partition of $\bm{x}$ into $M$ disjoint segments, which are ranked by their average attribution score. For $\kappa=0,\dots,K$, define the cumulative removal set $V_{\kappa}=\bigcup_{m=1}^{\kappa}S_{m}$ and perturbed input $\mathcal{P}_{\mathbb{X}}(\bm{x},V_{\kappa})$. The IROF metric $\Psi_{\mathrm{IROF}}$ is
@@
III.A.2 Limitations
@@
Faithfulness metrics inherit several limitations from perturbation-based explanation methods, as both rely on input perturbations to evaluate changes in the prediction certainty. Thus, they face similar issues to those in Section II.A.2, notably the need for carefully chosen baselines to avoid distribution shifts affecting model predictions. In addition, the FE metric relies on the assumption that removing a small number of highly attributed pixels will lead to a measurable change in model output. This assumption implicitly presumes that class-relevant information is concentrated in a small subset of pixels. However, RS scenes often contain multiple spatially disjoint regions associated with the same LULC class. These regions exhibit redundancy, as similar class-relevant features are repeated across the scene. In such cases, removing a small number of pixels does not produce a meaningful change in prediction, and the attribution sum remains nearly constant across different pixel subsets. As a result, the Pearson correlation used by FE may become numerically unstable, since both numerator and denominator can approach zero, leading to undefined or misleading scores even for faithful explanations.
In contrast to FE, trajectory-based metrics like IROF assess model behavior across progressively perturbed inputs. Their outcome heavily depends on the feature removal order. Let us consider a scene where LULC class $c_{\text{major}}$ covers most of the image and $c_{\text{minor}}$ only a small region. Perturbing $c_{\text{major}}$ pixels has little effect due to redundancy, so removing most relevant pixels first yields a slow decline in $f(\bm{x})$, as visualized in Fig. 4. In contrast, perturbing $c_{\text{minor}}$ pixels quickly lowers prediction certainty, biasing results toward faithful explanations for $c_{\text{minor}}$ and against $c_{\text{major}}$. The LeRF strategy, which removes the least relevant pixels first, mitigates this bias by preserving crucial features of $c_{\text{minor}}$ longer, stabilizing prediction certainty and producing more balanced evaluations.
@@
Figure 4: Comparison of MoRF and LeRF removal strategies for the target class $c$ (agricultural land, orange). Here, $K$ is the percentage of removed pixels and $f_{c}(\bm{x})$ represents the prediction certainty. Top row: a) Original image from the DG-ML [46] dataset; b) Pixel-wise reference map. Agriculture land in orange; MoRF removal: c) $K=10\%$, $f_{c}(\bm{x})=0.98$; d) $K=50\%$, $f_{c}(\bm{x})=0.64$; e) $K=90\%$, $f_{c}(\bm{x})=0.31$. Bottom row: f) Grad-CAM attribution map for $c$; LeRF removal: g) $K=10\%$, $f_{c}(\bm{x})=0.99$; h) $K=50\%$, $f_{c}(\bm{x})=0.99$; i) $K=90\%$, $f_{c}(\bm{x})=0.72$.
III.B Robustness
III.B.1 Background
@@
The second category we consider is robustness, which assesses the stability of explanations under small input perturbations. We require small perturbations $\delta$ such that $\left\|\delta\right\|_{2}<\varepsilon$ and expect $\left\|\mathcal{P}_{\mathbb{X}}(\bm{x})-\bm{x}\right\|_{2}<\varepsilon$ for a small positive constant $\varepsilon$, assuming that the model output approximately stayed the same $f(\bm{x})\approx f\left(\mathcal{P}_{\mathbb{X}}(\bm{x})\right)$. To evaluate the robustness of explanation methods for RS image scene classification, we use i) Average Sensitivity (AS) [30], which measures sensitivity to small perturbations using a Monte Carlo-based approximation and ii) Local Lipschitz Estimate (LLE) [28], which estimates Lipschitz continuity to assess explanation stability under minor input changes.
@@
AS [30] evaluates the sensitivity of the explanation method by perturbing samples using a Monte-Carlo-based approximation. The AS metric $\Psi_{\mathrm{AS}}$ is defined as:
@@
where $\varepsilon$ defines the radius of a discrete, finite-sample neighborhood around each input sample $\bm{x}$. This neighborhood, denoted as $\mathcal{N}_{\epsilon}(\bm{x})$, includes all samples that are within a distance of $\varepsilon$ from $\bm{x}$. A lower score is indicative of more robustness.
The LLE [28] estimates the Lipschitz continuity of the explanation, which measures how much the explanation changes with respect to the input under a minor perturbation, meaning the prediction did not change. The LLE metric $\Psi_{\mathrm{LLE}}$ is defined as:
@@
where, $\delta$ is the perturbation vector with $\delta\in\mathcal{N}_{\epsilon}(\bm{x})\leq\epsilon$, and $\mathcal{N}\varepsilon(\bm{x})$ is the neighborhood of perturbed samples within radius $\varepsilon$. Here, lower values indicate less change concerning the change in input.
III.B.2 Limitations
A key hyperparameter in this metric category is the noise type. If the introduced noise does not naturally occur in the acquired images, it may generate OoD samples, reducing the reliability of the model and thereby affecting the validity of the explanations. When evaluating the robustness of explanation methods, it is important to introduce noise that does not result in OoD samples. This consideration also depends on the model architecture, as many models exhibit inherent resilience to certain types of noise. For instance, in SAR images, multiplicative noise such as speckle is pervasive [54]. In contrast, for multispectral images, noise is typically independent of the data and generally additive in nature, such as additive white Gaussian noise [55, 56].
III.C Localization
III.C.1 Background
The third category we consider is localization, which measures the spatial accuracy of explanations by comparing them to reference maps or bounding boxes. This comparison can be quantified using a similarity measure SIM between the generated explanation map $\Phi(\bm{x})$ and a pixel-level reference map $\bm{s}_{gt}$. where SIM can be a measure such as Intersection over Union, Pearson correlation, or SSIM depending on the nature of the task and reference map. We utilize i) Top-K Intersection (TKI) [32], which measures the fraction of top-ranked attributions within the pixel-level reference map, and ii) Relevance Rank Accuracy (RRA) [31], which assesses the concentration of high-relevance values within the reference map.
The TKI metric [32] measures the fraction of the $K$ highest ranked attributions that lie on the binary reference mask $\bm{s}_{gt}$. We define $I_{K}=\{\,i_{1},i_{2},\ldots,i_{K}\}$ and $S_{gt}=\{\,p\mid\bm{s}_{gt}(p)=1\}$, where $I_{K}$ is the set of the top-$K$ ranked pixel indices and $S_{gt}$ the set of true-positive pixels in the mask. The intersection of these sets is $\bm{s}_{1-K}=I_{K}\;\cap\;S_{gt}$. The TKI metric $\Psi_{\mathrm{TKI}}$ is then defined as: